OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and third equations, then the second and third equations. x – y + 2z = –2 2x + 2y + z = 7 3x + 3y – z = 3

OpenStudy (anonymous):

2x + 2y = 7 3x + 3y = 3 2x + 2y = 7 5x + 5y = 10 7x + 5y = 4 3x + 3y = 3 7x + 5y = 4 5x + 5y = 10

OpenStudy (amorfide):

does using matrices count as an elimination method?

OpenStudy (campbell_st):

the last pair or equations are correct

OpenStudy (campbell_st):

5x + 5y = 10 is equation (2) + equation (3) 7x + 5y = 4 is equation (1) + 2 x equation (2)

OpenStudy (anonymous):

thank you